i must find the dimension of a rectangle a with the greatest area whose perimeter is 30 ft. is the answer 10 by 5
Answer by stanbon(75887)

i must find the dimension of a rectangle a with the greatest area whose perimeter is 30 ft. is the answer 10 by 5
---------------------
Area = length*width OR LW
-----------------------------
Perimeter = 2(L+W)
------
Your Problem:
30 = 2(L+W)
15 = L+W
W = 15-L
------------------
Area = LW
A = L(15-L)
A = 15L-L^2
------------
This is a quadratic equation with a = -1 and b = 15
The Maximum area occurs when L = -b/2a = -15/(-2) = 7.5
---------------------------------
Answer: Length = 7.5 ft and Width = 7.5 ft.
================================================
Cheers,
Stan H.